# Heavy Mesons to Charmed Tetraquark Decays

### v1

@@ VII Conclusion
  • We decompose $\overline{B}\to D\overline{T}$ and $\overline{B}\to\overline{D}T$ decay amplitudes in $\Delta S=-1$ and 00 transitions into topological amplitudes, where three topological amplitudes are needed in each case. However, only one topological amplitude, namely the $C_{\overline{T}}$ or $C_{\overline{D}T}$, is needed in $\overline{B}\to D\overline{T}$ or $\overline{B}\to\overline{D}T$ decays for modes involving flavor exotic $T$. Hence, their amplitudes are highly related. As there are many flavor exotic states in scenario II and the amplitudes of their production are highly related, the continued unobservation of these flavor exotic states in $\overline{B}\to D\overline{T}$ and $\overline{B}\to\overline{D}T$ decays poses tension in this scenario.
@@
  • A doubly charged tetraquark $T^{++}$ cannot be produced in $\overline{B}\to\overline{D}T$ decays, as such a decay violates charge conservation. In scenario I, one needs to consider both $\Delta S=-1$ and $\Delta S=0$ transitions in $\overline{B}\to D\overline{T}$ decays to have all four flavor exotic states participate. This is important as we need to verify whether $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$ are in the same multiplet or not. In scenario II, we also need to consider both $\Delta S=-1$ and $\Delta S=0$ transitions in $\overline{B}\to D\overline{T}$ decays. Note that the $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}\bar{s}}$ decay involving $\overline{T^{*}_{cs0}}(2870)^{0}$ has not been observed yet, as it is a CKM suppressed $\Delta S=-1$ transition. It will be interesting and useful to search for this mode to complete the above verification.
@@
  • The amplitudes of $\overline{B}\to TP$ and $\overline{B}\to TS$ decays in $\Delta S=0$ and $-1$ transitions are decomposed into topological amplitudes, where we need seven (seven) and eleven (nine) independent topological amplitudes, respectively, in scenario I (II). The number of topological amplitudes contributing to modes with flavor exotic states is reduced considerably. In $\overline{B}\to TP$ with flavor exotic $T$ and $P=\pi,K$, only three (four) independent combinations of topological amplitudes are needed in scenario I (II), resulting in several relations on rates of these modes. Similarly, in $\overline{B}\to TS$ with flavor exotic $T$ and $S=a_{0},\kappa$, only three (four) independent combinations of topological amplitudes are needed in scenario I (II), giving several rate relations. Note that from charge conservation, a doubly charged $T$ cannot be produced in $\overline{B}\to TP$ nor $\overline{B}\to TS$ decays.
@@
  • We note that $B^{-}_{c}\to T\overline{T}$ decays are kinematically allowed. The decay amplitudes can be decomposed into seven topological amplitudes, while only three of them contribute to modes with flavor exotic $T$ and $\overline{T}$. Interestingly, doubly charged $T$ can be produced in these decays. Furthermore, in scenario II, all flavor exotic states participate.
@@
  • A summary of the participation of flavor exotic states in scenario I in $\overline{B}\to D\overline{T}$, $\overline{B}\to\overline{T}$, $\overline{B}\to TP$, $\overline{B}\to TS$ and $B^{-}_{c}\to T\overline{T}$ decays in $\Delta S=0$ and $\Delta S=-1$ transitions is shown in Table 22, while the summary in scenario II is in Table 23.
@@ Appendix A SU(3) transformation of $q\bar{q}^{\prime}\bar{q}^{\prime\prime}$ states and SU(3) decompositions
$V_{i^{\prime}}{}^{i^{\prime}k}=({\cal T}_{a})^{i}_{i^{\prime}}{}_{j}^{i^{\prime}}{}_{k}^{k^{\prime}}V_{i}{}^{jk}=0,\quad V_{i^{\prime}}{}^{k^{\prime}i^{\prime}}=({\cal T}_{a})^{i}_{i^{\prime}}{}_{j}^{j^{\prime}}{}_{k}^{i^{\prime}}V_{i}{}^{jk}=0.$ (81)
Consequently, SU(3) transformation does not mix up $\bar{q}^{\prime}\bar{q}^{\prime\prime}$ in different permutation symmetry, nor does it mix up traceless $q\bar{q}^{\prime}\bar{q}^{\prime\prime}$ combinations with the traceful combinations. Therefore, we can decompose $\bar{q}^{\prime}\bar{q}^{\prime\prime}$ into $[\bar{q}^{\prime}\bar{q}^{\prime\prime}]$ and $\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}$ at the outset. Hence, we have scenario I and scenario II, where scenario I has $T=T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]}$, while scenario II has $T=T_{cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}}$, and they can be further separated into traceless and traceful parts.
@@
and their hermitian conjugations.
@@
One can easily check using the above generators that the states in ${\bf 6}$ and $\bar{\bf 3}$ in scenario I, as shown in Table 1, and states in $\overline{\bf 15}$ and $\bar{\bf 3}^{\prime}$ in scenario II, as shown Table 2, are in their own irreducible representations after SU(3) transformation.
Furthermore, the isospin multiplets within those irreducible representations can be verified using ${\cal I}={\cal T}_{3}$, ±${\cal I}_{\pm}={\bf\cal T}_{1}\pm i{\bf\cal T}_{2}$ and ${\cal I}^{2}$ with
${\cal I}^{2}=({\cal T}_{3})^{2}+\frac{1}{2}({\cal T}_{1}+i{\cal T}_{2})({\cal T}_{1}-i{\cal T}_{2})+\frac{1}{2}({\cal T}_{1}-i{\cal T}_{2})({\cal T}_{1}+i{\cal T}_{2}).$ (132)
The results are shown in Tables 1 and 2.
@@
$\bar{\bf 3}\otimes\bar{\bf 3}=\bar{\bf 6}\oplus{\bf 3},\quad{\bf 3}\otimes\bar{\bf 6}=\overline{\bf 15}\oplus\bar{\bf 3},\quad{\bf 3}\otimes{\bf 3}={\bf 6}\oplus\bar{\bf 3},$ (133)
$\bar{\bf 3}\otimes{\bf 8}$ $=$ $\overline{\bf 15}\oplus{\bf 6}\oplus\bar{\bf 3},\quad\bar{\bf 3}\otimes{\bf 1}=\bar{\bf 3},$

### v3

@@ VII Conclusion
  • We decompose $\overline{B}\to D\overline{T}$ and $\overline{B}\to\overline{D}T$ decay amplitudes in $\Delta S=-1$ and 0 transitions into topological amplitudes, where six topological amplitudes are needed in each case. However, only $C_{\overline{T}}+P_{\bar{T},1}$ or $C_{\overline{D}T}+P_{T}$, are needed in $\overline{B}\to D\overline{T}$ or $\overline{B}\to\overline{D}T$ decays for modes involving flavor exotic $T$. Hence, their amplitudes are highly related. As there are many flavor exotic states in scenario II and the amplitudes of their production are highly related, the continued unobservation of these flavor exotic states in $\overline{B}\to D\overline{T}$ and $\overline{B}\to\overline{D}T$ decays poses tension in this scenario.
@@
  • A doubly charged open-charmed tetraquark $T^{++}$ cannot be produced in $\overline{B}\to\overline{D}T$ decays, as such a decay violates charge conservation. In scenario I, one needs to consider both $\Delta S=-1$ and $\Delta S=0$ transitions in $\overline{B}\to D\overline{T}$ decays to have all four flavor exotic states participate. This is important as we need to verify whether $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$ are in the same multiplet or not. In scenario II, we also need to consider both $\Delta S=-1$ and $\Delta S=0$ transitions in $\overline{B}\to D\overline{T}$ decays. Note that the $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}\bar{s}}$ decay involving $\overline{T^{*}_{cs0}}(2870)^{0}$ has not been observed yet, as it is a CKM suppressed $\Delta S=-1$ transition. It will be interesting and useful to search for this mode to complete the above verification. It should be noted that there is penguin pollution in the above programs. Although the effects on rate ratios are estimated to be mild, further works are needed for verification. We also note that the $\overline{B}^{0}\to\overline{D}^{0}T^{0}_{cs}$ decay has not been observed yet. It will be useful to search for this mode and check the relations on rates.
@@
  • The amplitudes of $\overline{B}\to TP$ and $\overline{B}\to TS$ decays in $\Delta S=0$ and 1-1 transitions are decomposed into topological amplitudes, where we need seven (seven) and eleven (nine) independent topological amplitudes, respectively, in scenario I (II). The number of topological amplitudes contributing to modes with flavor exotic states is reduced considerably. In $\overline{B}\to TP$ with flavor exotic $T$ and $P=\pi,K$, only three (four) independent combinations of topological amplitudes are needed in scenario I (II), resulting in several relations on rates of these modes. Similarly, in $\overline{B}\to TS$ with flavor exotic $T$ and $S=a_{0},\kappa$, only three (four) independent combinations of topological amplitudes are needed in scenario I (II), giving several rate relations. Although the doubly charged $T^{*}_{c\bar{s}0}(2900)^{++}$ cannot be produced in $\overline{B}\to TP$ nor $\overline{B}\to TS$ decays, there are relations that can be used to verify whether $T_{cs0}(2870)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{0}$ are in the same multiplets or not. According to the OZI rule, the $\overline{B}\to TS$ topological amplitudes are expected to receive some suppressions.
@@
  • We note that $B^{-}_{c}\to T\overline{T}$ decays are kinematically allowed. Interestingly, doubly charged $T$ can be produced in these decays. The $B^{-}_{c}\to T\overline{T}$ decay amplitudes can be decomposed into seven topological amplitudes, while only three of them contribute to modes with flavor exotic $T$ and $\overline{T}$. In scenario I, these three topological amplitudes only occur in one combination for modes with flavor exotic $T$ and $\overline{T}$ ; these amplitudes are all related. In scenario II, all flavor exotic states participate in these decays. According to the OZI rule, the topological amplitudes in $B_{c}^{-}\to T\overline{T}$ decays are expected to have some suppressions.
@@
  • A summary of the participation of flavor exotic states in scenario I in $\overline{B}\to D\overline{T}$, $\overline{B}\to\overline{T}$, $\overline{B}\to TP$, $\overline{B}\to TS$ and $B^{-}_{c}\to T\overline{T}$ decays in $\Delta S=0$ and $\Delta S=-1$ transitions is shown in Table 22, while the summary for scenario II is in Table 23.
@@ Appendix A SU(3) transformation of $q\bar{q}^{\prime}\bar{q}^{\prime\prime}$ states and SU(3) decompositions
$V_{i^{\prime}}{}^{i^{\prime}k}=({\cal T}_{a})^{i}_{i^{\prime}}{}_{j}^{i^{\prime}}{}_{k}^{k^{\prime}}V_{i}{}^{jk}=0,\qquad V_{i^{\prime}}{}^{k^{\prime}i^{\prime}}=({\cal T}_{a})^{i}_{i^{\prime}}{}_{j}^{j^{\prime}}{}_{k}^{i^{\prime}}V_{i}{}^{jk}=0.$ (112)
Consequently, SU(3) transformation does not mix up $\bar{q}^{\prime}\bar{q}^{\prime\prime}$ in different permutation symmetry, nor does it mix up traceless $q\bar{q}^{\prime}\bar{q}^{\prime\prime}$ combinations with the traceful combinations. Therefore, we can decompose $\bar{q}^{\prime}\bar{q}^{\prime\prime}$ into $\bar{q}^{\prime}\bar{q}^{\prime\prime}]$ and $\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}$ at the outset. Hence, we have scenario I and scenario II, where scenario I has $T=T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]}$, while scenario II has $T=T_{cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}}$, and they can be further separated into traceless and traceful parts.
@@
Table 24: The results of SU(3) generators acting on states in scenario I with quark contents given in Table 1. They are in their own irreducible representations after these transformations. Comparing this table with Fig. 1, one can verify that these transformations indeed follow Eq. (173).
$T_{c}$ ${\cal I}_{3}$ $\cal S$} ${\cal T}_{1}+i{\cal T}_{2}$ ${\cal T}_{1}-i{\cal T}_{2}$ ${\cal T}_{4}+i{\cal T}_{5}$ ${\cal T}_{4}-i{\cal T}_{5}$ ${\cal T}_{6}+i{\cal T}_{7}$ ${\cal T}_{6}-i{\cal T}_{7}$
$T^{++}_{c\bar{s}} +$$T^{++}_{c\bar{s}} +$$T^{++}_{c\bar{s}}$ 0 $\sqrt{2}T^{+}_{c\bar{s}}$ 0 $\sqrt{2}T_{c}^{+}$ 0 0
$T^{+}_{c\bar{s}}$ 0 $T^{+}_{c\bar{s}}$ $\sqrt{2}T^{++}_{c\bar{s}}$ $\sqrt{2}T^{0}_{c\bar{s}}$ 0 $T^{0}_{c}$ 0 $T^{+}_{c}$
$T^{0}_{c\bar{s}}$ $T^{0}_{c\bar{s}}$ $T^{0}_{c\bar{s}}$ $\sqrt{2}T_{c\bar{s}}^{+}$ 0 0 0 0 $\sqrt{2}T^{0}_{c}$
$T^{0}_{cs}$ 0 $T^{0}_{cs}$ 0 0 $\sqrt{2}T^{+}_{c}$ 0 $\sqrt{2}T^{0}_{c}$ 0
$T^{+}_{c} +$$\frac{1}{2}T^{+}_{c}$ 0 0 $T_{c}^{0}$ $\sqrt{2}T^{++}_{c\bar{s}}$ $\sqrt{2}T^{0}_{cs}$ $T_{c\bar{s}}^{+}$ 0
$T^{0}_{c}$ $\frac{1}{2}T^{0}_{c}$ 0 $T^{+}_{c}$ 0 $T^{+}_{c\bar{s}}$ 0 $\sqrt{2}T_{c\bar{s}}^{0}$ $\sqrt{2}T_{cs}^{0}$
$T^{\prime\prime+}_{c} +$$\frac{1}{2}T^{\prime\prime+}_{c}$ 0 0 $T^{\prime\prime 0}_{c}$ 0 0 $T^{\prime\prime+}_{c\bar{s}}$ 0
$T^{\prime\prime 0}_{c}$ $\frac{1}{2}T^{\prime\prime 0}_{c}$ 0 $T^{\prime\prime+}_{c}$ 0 $T^{\prime\prime+}_{c\bar{s}}$ 0 0 0
$T^{\prime\prime+}_{c\bar{s}}$ 0 $T^{\prime\prime+}_{c\bar{s}}$ 0 0 0 $T^{\prime\prime 0}_{c}$ 0 $T^{\prime\prime+}_{c}$
@@
and their hermitian conjugations, while generators for isospin ${\cal I}_{3}$ and stangeness $\cal S$} can be defined as
${\cal I}$ $=$ ${\cal T}_{3},$
${\cal S}$ $\equiv$ ${\mathbb{I}}\otimes{\mathbb{I}}\otimes{\mathbb{I}}+2\sqrt{3}{\cal T}_{8}$ (172)
$=$ $-{\mathbb{I}}\otimes{\mathbb{I}}\otimes{\mathbb{I}}+{\mathbb{I}}\otimes{\mathbb{I}}\otimes{\mathbb{I}}+{\mathbb{I}}\otimes{\mathbb{I}}\otimes{\mathbb{I}}+2\sqrt{3}{\cal T}_{8}$
$=$ $\left(\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&-1\\ \end{array}\right)\otimes{\mathbb{I}}\otimes{\mathbb{I}}-{\mathbb{I}}\otimes\left(\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&-1\\ \end{array}\right)\otimes{\mathbb{I}}-{\mathbb{I}}\otimes{\mathbb{I}}\otimes\left(\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&-1\\ \end{array}\right).$
These generators are responsible for generating SU(3) transformation. In particular, they will move states with quantum numbers $(I_{3},S)$ in an irreducible multiplet to other states in the same multiplet in the following way:
${\cal T}_{1}\pm i{\cal T}_{2}$ $:$ $(I_{3},S)\to(I_{3}\pm 1,S),$
${\cal T}_{4}\pm i{\cal T}_{5}$ $:$ $(I_{3},S)\to(I_{3}\pm 1,S\pm 1),$
${\cal T}_{6}\pm i{\cal T}_{7}$ $:$ $(I_{3},S)\to(I_{3}\mp 1,S\pm 1).$ (173)
These generators are known as the raising and lowering operators of isospin, $V$-spin and $U$-spin, respectively. Note that if the transformed state does not exist in the multiplet, the result of that transformation will end up in a null state.
Table 25: Same as Table. 24 but for scenario II with states defined in Table 2.
$T_{c}$ ${\cal I}_{3}$ $\cal S$} ${\cal T}_{1}+i{\cal T}_{2}$ ${\cal T}_{1}-i{\cal T}_{2}$ ${\cal T}_{4}+i{\cal T}_{5}$ ${\cal T}_{4}-i{\cal T}_{5}$ ${\cal T}_{6}+i{\cal T}_{7}$ ${\cal T}_{6}-i{\cal T}_{7}$
$T^{++}_{c\bar{s}} +$$T^{++}_{c\bar{s}} +$$T^{++}_{c\bar{s}}$ 0 $\sqrt{2}T^{+}_{c\bar{s}}$ 0 $\frac{\sqrt{2}T_{c}^{+}+2T^{\prime+}_{c}}{\sqrt{3}}$ $\sqrt{2}T^{++}_{c\bar{s}\bar{s}}$ $\sqrt{2}T_{c}^{++}$
$T^{+}_{c\bar{s}}$ 0 $T^{+}_{c\bar{s}}$ $\sqrt{2}T^{++}_{c\bar{s}}$ $\sqrt{2}T^{0}_{c\bar{s}}$ $T^{++}_{c\bar{s}\bar{s}}$ $\frac{2T^{0}_{c}+\sqrt{2}T^{\prime 0}_{c}}{\sqrt{3}}$ $T_{c\bar{s}\bar{s}}^{+}$ $\frac{-2T^{+}_{c}+\sqrt{2}T^{\prime+}_{c}}{\sqrt{3}}$
$T^{0}_{c\bar{s}}$ $T^{0}_{c\bar{s}}$ $T^{0}_{c\bar{s}}$ $\sqrt{2}T_{c\bar{s}}^{+}$ 0 $\sqrt{2}T_{c\bar{s}\bar{s}}^{+}$ $\sqrt{2}T_{c}^{-}$ 0 $\frac{\sqrt{2}T^{0}_{c}-2T^{\prime 0}_{c}}{\sqrt{3}}$
$T^{+}_{cs} +$$T^{+}_{cs}$ $T^{+}_{cs}$ 0 $\sqrt{2}T^{0}_{cs}$ $T^{++}_{c}$ 0 $\frac{-T^{+}_{c}+2\sqrt{2}T^{\prime+}_{c}}{\sqrt{3}}$ 0
$T^{0}_{cs}$ 0 $T^{0}_{cs}$ $\sqrt{2}T^{+}_{cs}$ $\sqrt{2}T^{-}_{cs}$ $\frac{\sqrt{2}T^{+}_{c}+2T^{\prime+}_{c}}{\sqrt{3}}$ 0 $\frac{-\sqrt{2}T^{0}_{c}+2T^{\prime 0}_{c}}{\sqrt{3}}$ 0
$T^{-}_{cs}$ $T^{-}_{cs}$ $T^{-}_{cs}$ $\sqrt{2}T^{0}_{cs}$ 0 $\frac{T^{0}_{c}+2\sqrt{2}T^{\prime 0}_{c}}{\sqrt{3}}$ 0 $T^{-}_{c}$ 0
$T^{++}_{c\bar{s}\bar{s}} +$$\frac{1}{2}T^{++}_{c\bar{s}\bar{s}} +$$2T^{++}_{c\bar{s}\bar{s}}$ 0 $T^{+}_{c\bar{s}\bar{s}}$ 0 $T^{+}_{c\bar{s}}-\sqrt{2}T^{\prime+}_{c\bar{s}}$ 0 $\sqrt{2}T^{++}_{c\bar{s}}$
$T^{+}_{c\bar{s}\bar{s}}$ $\frac{1}{2}T^{+}_{c\bar{s}\bar{s}} +$$2T^{+}_{c\bar{s}\bar{s}}$ $T^{++}_{c\bar{s}\bar{s}}$ 0 0 $\sqrt{2}T^{0}_{c\bar{s}}$ 0 $T^{+}_{c\bar{s}}-\sqrt{2}T^{\prime+}_{c\bar{s}}$
$T^{++}_{c} +$$\frac{3}{2}T^{++}_{c}$ 0 0 $\sqrt{3}T^{+}_{c}$ 0 $T^{+}_{cs}$ $\sqrt{2}T^{++}_{c\bar{s}}$ 0
$T^{+}_{c} +$$\frac{1}{2}T^{+}_{c}$ 0 $\sqrt{3}T^{++}_{c}$ $2T_{c}^{0}$ $\sqrt{\frac{2}{3}}T^{++}_{c\bar{s}}$ $\sqrt{\frac{2}{3}}T^{0}_{cs}$ $\frac{2T_{c\bar{s}}^{+}}{\sqrt{3}}$ $\frac{T_{cs}^{+}}{\sqrt{3}}$
$T^{0}_{c}$ $\frac{1}{2}T^{0}_{c}$ 0 $2T^{+}_{c}$ $\sqrt{3}T_{c}^{-}$ $\frac{2T^{+}_{c\bar{s}}}{\sqrt{3}}$ $\frac{T^{-}_{cs}}{\sqrt{3}}$ $\sqrt{\frac{2}{3}}T_{c\bar{s}}^{0}$ $\sqrt{\frac{2}{3}}T_{cs}^{0}$
$T^{-}_{c}$ $\frac{3}{2}T^{-}_{c}$ 0 $\sqrt{3}T^{0}_{c}$ 0 $\sqrt{2}T^{0}_{c\bar{s}}$ 0 0 $T^{-}_{cs}$
$T^{\prime+}_{c} +$$\frac{1}{2}T^{\prime+}_{c}$ 0 0 $T^{\prime 0}_{c}$ $\frac{2T_{c\bar{s}}^{++}}{\sqrt{3}}$ $\frac{2T_{cs}^{0}}{\sqrt{3}}$ $\frac{\sqrt{2}T_{c\bar{s}}^{+}-3T^{\prime+}_{c\bar{s}}}{\sqrt{3}}$ $2\sqrt{\frac{2}{3}}T_{cs}^{+}$
$T^{\prime 0}_{c}$ $\frac{1}{2}T^{\prime 0}_{c}$ 0 $T^{\prime+}_{c}$ 0 $\frac{\sqrt{2}T_{c\bar{s}}^{+}+3T^{\prime+}_{c\bar{s}}}{\sqrt{3}}$ $2\sqrt{\frac{2}{3}}T_{cs}^{-}$ $\frac{2T_{c\bar{s}}^{0}}{\sqrt{3}}$ $\frac{2T_{cs}^{0}}{\sqrt{3}}$
$T^{\prime+}_{c\bar{s}}$ 0 $T^{\prime+}_{c\bar{s}}$ 0 0 $\sqrt{2}T_{c\bar{s}\bar{s}}^{++}$ $\sqrt{3}T^{\prime 0}_{c}$ $\sqrt{2}T_{c\bar{s}\bar{s}}^{+}$ $\sqrt{3}T^{\prime+}_{c}$
$T^{\prime\prime+}_{c} +$$\frac{1}{2}T^{\prime\prime+}_{c}$ 0 0 $T^{\prime\prime 0}_{c}$ 0 0 $T^{\prime\prime+}_{c\bar{s}}$ 0
$T^{\prime\prime 0}_{c}$ $\frac{1}{2}T^{\prime\prime 0}_{c}$ 0 $T^{\prime\prime+}_{c}$ 0 $T^{\prime\prime+}_{c\bar{s}}$ 0 0 0
$T^{\prime\prime+}_{c\bar{s}}$ 0 $T^{\prime\prime+}_{c\bar{s}}$ 0 0 0 $T^{\prime\prime 0}_{c}$ 0 $T^{\prime\prime+}_{c}$
@@
As shown in Tables 24 and 25, using the above generators the states in $\bf 6$} and $\bar{\bf 3}$ in scenario I, given in Table 1, and states in $\overline{\bf 15}$ and $\bar{\bf 3}^{\prime}$ in scenario II, given in Table 2, are in their own irreducible representations after SU(3) transformations. By comparing Tables 24 and 25 and Fig. 1, one can also verify that these transformations indeed follow Eq. (173 ). Furthermore, the isospin multiplets within those irreducible representations can be verified using ${\cal I}_{\pm}={\bf\cal T}_{1}\pm i{\bf\cal T}_{2}$ and ${\cal I}^{2}$ with
${\cal I}^{2}=({\cal I}_{3})^{2}+\frac{1}{2}({\cal T}_{1}+i{\cal T}_{2})({\cal T}_{1}-i{\cal T}_{2})+\frac{1}{2}({\cal T}_{1}-i{\cal T}_{2})({\cal T}_{1}+i{\cal T}_{2}).$ (174)
This quantum number can be obtained easily using the results in Tables 24 and 25. The final results are shown in Tables 1 and 2.
@@
$\bar{\bf 3}\otimes\bar{\bf 3}=\bar{\bf 6}\oplus{\bf 3},\qquad{\bf 3}\otimes\bar{\bf 6}=\overline{\bf 15}\oplus\bar{\bf 3},\qquad{\bf 3}\otimes{\bf 3}={\bf 6}\oplus\bar{\bf 3},$ (175)
$\bar{\bf 3}\otimes{\bf 8}$ $=$ $\overline{\bf 15}\oplus{\bf 6}\oplus\bar{\bf 3},\qquad\bar{\bf 3}\otimes{\bf 1}=\bar{\bf 3},$